25. An analysis of variance comparing 3 treatment conditions produces dfwithin 21. If the samples are all the same size, how many individuals are in each sample? a, 6 b. 7 c. 8 d. it is impossible for the samples to be the same size if dfwithin 21
Answer #1

here treatment(k) =3 and n= 21

and we know that n=N-k

where n is dfwithin N is total degree of freedom

so N= 24

indivodaula are in each sample is = N/k

= 24÷3 =8

hence (c) 8 is right answer